OpenStudy (anonymous):

|2n - 3| = 5 A. {-1, 4} B. {1, 4} C. {-1, 1} D. {1, -4}

OpenStudy (ajprincess):

|2n - 3| = 5 2n-3=5 or 2n-3=-5 Solve for n. Does that help?

OpenStudy (ajprincess):

@Zoobidoo

OpenStudy (anonymous):

Oh, hi. It sort of helps.

OpenStudy (anonymous):

Im not sure what steps follow though...

OpenStudy (nurali):

|2n - 3| = 5 2n-3=5 or 2n-3=-5 2n=5+3 or 2n=-5+3 2n=8 or 2n=-2 n=8/2 or n=-2/2 n=4 or n=-1

OpenStudy (ajprincess):

2n-3=5 1.Add 3 both sides. 2.Divide both sides by 2 Pls follow the above two steps to solve for n.

OpenStudy (anonymous):

a -5=2n-3 / 2n-3=5 ==> -2=2n , 2n =8 n=4,n=-1

OpenStudy (anonymous):

Ok I get it! This helps so much! I really appreciate it!

OpenStudy (anonymous):

@nurali @mahdi2020 @ajprincess Thank you!

OpenStudy (ajprincess):

Welcome:)

OpenStudy (nurali):

Anytime.